# The Thinking Therapist: Training Large Language Models to Deliver Acceptance and Commitment Therapy using Supervised Fine-Tuning and Odds Ratio Policy Optimization

### v1

@@ Abstract
Acceptance and Commitment Therapy (ACT) is a third-wave cognitive behavioral therapy with emerging evidence of efficacy in several psychiatric conditions. This study investigates the impact of post-training methodology and explicit reasoning on the ability of a small open-weight large language model (LLM) to deliver ACT. Using 50 sets of synthetic ACT transcripts generated by Mistral-Large, we trained Llama-3.2-3b-Instruct with two distinct approaches, supervised fine-tuning (SFT) and odds ratio policy optimization (ORPO), each with and without an explicit chain-of-thought (COT) reasoning step. Performance was evaluated by comparing these four post-trained variants against the base Instruct model. These models were benchmarked in simulated therapy sessions, with performance quantitatively assessed on the ACT Fidelity Measure (ACT-FM) and the Therapist Empathy Scale (TES) by an LLM judge that had been fine-tuned on human evaluations. Our findings demonstrate that the ORPO-trained models significantly outperformed both their SFT and Instruct counterparts on ACT fidelity ($\chi^{2}(5)=185.15,p<.001$) and therapeutic empathy ($\chi^{2}(5)=140.37,p<.001$). The effect of COT was conditional as it provided a significant benefit to SFT models, improving ACT-FM scores by an average of 2.68 points ($p<.001$), while offering no discernible advantage to the superior ORPO or instruct-tuned variants. We posit that the superiority of ORPO stems from its ability to learn the therapeutic ‘process’ over imitating ‘content,’ a key aspect of ACT, while COT acts as a necessary scaffold for models trained only via imitation. This study establishes that preference-aligned policy optimization can effectively instill ACT competencies in small LLMs, and that the utility of explicit reasoning is highly dependent on the underlying training paradigm.
Keywords Large Language Models  $\cdot$ Psychotherapy  $\cdot$ AI psychotherapy  $\cdot$ Acceptance and Commitment Therapy  $\cdot$ Reinforcement Learning  $\cdot$ Preference Optimization  $\cdot$ Chain-of-Thought
@@ 2.2 Synthetic Data Generation
The patient profiles were then used in the production of 100 synthetic ACT transcripts. Each transcript contained 25 turns of conversation. Mistral-Large was employed to generate both patient and therapist responses. Therapist responses were structured to have the thinking tokens appear after the <thinking> tag and final response tokens appear after the <answer> tag.
For quality assurance, DeepSeek-Chat was used as an LLM supervisor. Patient responses were inspected for consistency and realism whereas therapist responses were inspected for structure, ACT fidelity, context appropriateness and coherence. The LLM supervisor provided a binary yes or no verdict as to whether revision was required and offered textual feedback for improvement. Responses that required regeneration were sent back to Mistral-Large with the feedback from the LLM supervisor. Employing DeepSeek-Chat as the supervisor provided an orthogonal assessment of quality, to reduce model-specific artifacts and prevent the compounding of the generator’s inductive biases.
@@ 2.4 Model Testing and Evaluation
The performance of the trained models was evaluated through interactions with simulated patients. The ‘patient agent’ in the simulations was gpt-5-mini and the ‘therapist agent’ was one of six possible models and configurations, instruct-tuned (as released by Meta), instruct-tuned with a prompt that mandated producing responses in a Thinking:/Answer: format, fine-tuned on the synthetic ACT dataset with and without COT and ORPO trained again with and without COT. The goal was to be able to compare the performance of these six distinct approaches in delivering ACT-informed responses, thereby isolating the impact of supervised fine-tuning, chain-of-thought reasoning, and reinforcement learning on therapeutic quality.
@@ 2.5.3 Data Availability
All python code used for the primary analysis and exploratory subgroup analyses, along with the synthetic training data and evaluation results, are available in the project’s GitHub repository. The SFT and ORPO models are available on HuggingFace/TTahir.
@@ 3.2 Analysis of Therapeutic Quality
(a) ACT-FM Total Score
(b) TES Mean Score
Figure 1: Distribution of primary outcome scores across the six model variants. The central line indicates the median, the box represents the interquartile range (IQR), and the whiskers extend to 1.5 times the IQR.
@@ 3.2.2 Therapeutic Empathy
(a) Pairwise Comparisons for ACT-FM
(b) Pairwise Comparisons for TES
Figure 2: Forest plots illustrating the estimated mean differences and 95% confidence intervals from post-hoc pairwise comparisons for both primary outcomes. All comparisons are relative to the model listed on the y-axis. A positive difference indicates the model on the x-axis scored higher.
@@ 3.4.2 Patient Specific Performance
Figure 3: Heatmap of Mean ACT-FM total consistency scores by model variant and patient primary issue. Deeper green shades indicate better performance, while red shades indicate lower scores.
@@ 4.5 Strengths and Limitations
Our study was also limited by the small size of the training data and base model. It is entirely possible that scaling laws could shift the relative importance of COT and RL at larger capacities.

### v2

@@ Abstract
Acceptance and Commitment Therapy (ACT) is a third-wave cognitive behavioral therapy with emerging evidence of efficacy in several psychiatric conditions. This study investigates the impact of post-training methodology and explicit reasoning on the ability of a small open-weight large language model (LLM) to deliver ACT. Using synthetic ACT transcripts generated by Mistral-Large, we trained Llama-3.2-3b-Instruct with two distinct approaches, supervised fine-tuning (SFT) and odds ratio policy optimization (ORPO), each with and without an explicit chain-of-thought (COT) reasoning step. Performance was evaluated by comparing these four post-trained variants against the base Instruct model. These models were benchmarked in simulated therapy sessions, with performance quantitatively assessed on the ACT Fidelity Measure (ACT-FM) and the Therapist Empathy Scale (TES) by an LLM judge that had been fine-tuned on human evaluations. Our findings demonstrate that the ORPO-trained models significantly outperformed both their SFT and Instruct counterparts on ACT fidelity ($\chi^{2}(5)=185.15,p<.001$) and therapeutic empathy ($\chi^{2}(5)=140.37,p<.001$). The effect of COT was conditional as it provided a significant benefit to SFT models, improving ACT-FM scores by an average of 2.68 points ($p<.001$), while offering no discernible advantage to the superior ORPO or instruct-tuned variants. We posit that the superiority of ORPO stems from its ability to learn the therapeutic ‘process’ over imitating ‘content,’ a key aspect of ACT, while COT acts as a necessary scaffold for models trained only via imitation. This study establishes that preference-aligned policy optimization can effectively instill ACT competencies in small LLMs, and that the utility of explicit reasoning is highly dependent on the underlying training paradigm.
@@ 2.2 Synthetic Data Generation
The patient profiles were then used in the production of 100 synthetic ACT transcripts. Each transcript contained 25 turns of conversation. Mistral-Large was employed to generate both patient and therapist responses. Therapist responses were structured to have the reasonsing tokens appear after the <thinking> tag and final response tokens appear after the <answer> tag.
For quality assurance, DeepSeek-Chat was used as an LLM supervisor. Patient responses were inspected for consistency and realism whereas therapist responses were inspected for structure, ACT fidelity, context appropriateness and coherence. The LLM supervisor provided a binary ‘yes’ or ‘no’ verdict as to whether revision was required and offered textual feedback for improvement. Responses that required regeneration were sent back to Mistral-Large with the feedback from the LLM supervisor. Employing DeepSeek-Chat as the supervisor provided an orthogonal assessment of quality, to reduce model-specific artifacts and prevent compounding of the generator’s inductive biases.
@@ 2.4 Model Testing and Evaluation
The performance of the trained models was evaluated through interactions with simulated patients. The ‘patient agent’ in the simulations was gpt-5-mini and the ‘therapist agent’ was one of six possible models and configurations, instruct-tuned (as released by Meta), instruct-tuned with a prompt that mandated producing responses in a Thinking:/Answer: format, fine-tuned on the synthetic ACT dataset with and without COT, and ORPO trained again with and without COT. The goal was to be able to compare the performance of these six distinct approaches in delivering ACT-informed responses, thereby isolating the impact of supervised fine-tuning, chain-of-thought reasoning, and reinforcement learning on therapeutic quality.
@@ 2.5.3 Data Availability
All python code used for the primary analysis and exploratory subgroup analyses, along with the synthetic training data and evaluation results, can be accessed through the project’s GitHub repository. The SFT and ORPO models are available at HuggingFace/TTahir.
@@ 3.2 Analysis of Therapeutic Quality
Figure 1: Model Performance on Primary Outcomes. Distribution of primary outcome scores for ACT-FM and TES across the six model variants. The central line indicates the median, the box represents the interquartile range (IQR), and the whiskers extend to 1.5 times the IQR.
@@ 3.2.2 Therapeutic Empathy
Figure 2: Pairwise Model Comparisons on Therapeutic Quality. Forest plots illustrating estimated mean differences and 95% confidence intervals from post-hoc pairwise comparisons for ACT-FM and TES. All comparisons are relative to the model listed on the y-axis. A positive difference indicates the model on the x-axis scored higher.
@@ 3.4.2 Patient Specific Performance
Figure 3: Model Performance Across Patient Clinical Presentations. Heatmap of Mean ACT-FM total consistency scores by model variant and patient primary issue. Deeper green shades indicate better performance, while red shades indicate lower scores.
@@ 4.5 Strengths and Limitations
Our study was also limited by the small size of the training data and base model. It is entirely possible that scaling laws could shift the relative importance of COT and reinforcement learning at larger capacities.